# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ 3.3 Prior elicitation
We now focus on the concentration parameter of the DP prior in (5). It is clear from the posterior in (6) that $c=0$ corresponds to the prior belief that there is no ME present on the covariates resulting in all the mass of the posterior centering measure concentrating at $\delta_{w_{i}}$. This is in contrast to the standard NPL setting (Sec. 2) where $c=0$ corresponds to a non-informative prior. Moreover, it is evident that $c=1$ results in splitting the weight of the DP posterior centering measure $\operatorname{\mathbb{F}}_{w_{i}}^{\prime}$ equally between the prior centering measure $\operatorname{\mathbb{F}}_{w_{i}}$ and the Dirac measure on the observation $w_{i}$. This choice splits the information between the prior and observation. Lastly, when $c\rightarrow\infty$ we see that all the weight is placed in the prior belief.
The interplay between parameter $c$, ME variance and prior specification can be further demonstrated with an example of polynomial regression with Berkson, Normally distributed ME in Figure 3. Here, the average MSE over the three parameters of interest is reported for the three limiting choices of $c$, for an increasing ME variance. In the case of a well-specified prior (left), when the ME variance is close to zero, i.e. when there is no ME, all three $c$ values perform similarly since the prior correctly indicates no ME, making inference data-driven. Choices of $c>0$ perform similarly, yielding lower MSE for non-zero ME variance. Conversely, in the case where the true ME distribution is Gaussian and the prior centering measure $\operatorname{\mathbb{F}}_{N}^{0}$ is a Student-t distribution with 3 degrees of freedom (right plot), $c=1$ corresponds to the lowest average MSE, as expected. A setting of $c=1$ appropriately balances prior beliefs and observations, whereas $c=0$ wrongly assumes no ME, and $c=100$ overly weights the misspecified prior. Overall, overlapping standard deviations in most cases indicate the method’s robustness to $c$ values. In the absence of strong prior beliefs on the ME distribution, setting $c=1$ splits the weight sensibly between prior beliefs and observations.
Figure 3: Average MSE of polynomial regression parameters in the presence of Berkson, Gaussian ME with increasing standard deviation over 50 replications. Left: Robust-MEM method with well-specified prior results in a lower MSE for non-zero values of $c$. Right: Robust-MEM method with a misspecified (Student-t with 3 degrees of freedom) prior. The case of $c=1$ which splits the posterior weight between the data and the prior results to the lowest MSE in most cases.
@@ 3.4 Choice of loss function
In this section, we explore the adaptability of the Robust-MEM framework to various prevalent cases of ME. We propose two distinct loss functions suitable to address these scenarios comprehensively. The initial choice stems from the Total Least Squares objective, offering rapid and robust solutions, particularly tailored for scenarios such as linear regression and Gaussian-distributed ME. To ensure broad applicability across diverse model families and regression relationships, we introduce the MMD-based loss function as a further approach.
@@ 3.4.1 Robust-MEM with the Total Least Squares
We first suggest a loss function based on the TLS objective function. Consider the linear regression model in (2). Then we can define the loss function $l_{\operatorname{\operatorname{TLS}}}:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow\mathbb{R}$ as
$l_{\operatorname{\operatorname{TLS}}}(x_{i},y_{i},\theta):=\|\nu_{i}\|_{2}^{2}+(y_{i}-\theta^{T}(x_{i}+\nu_{i}))^{2}.$
Using this loss function in (8), the minimization objective for each set of DP samples $(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ as in 10 is
$\theta^{*}_{TLS}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ $=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{(x,y)\sim\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\delta_{y_{i}}}\left[l_{\operatorname{\operatorname{TLS}}}(x,y,\theta)\right]$ (12)
$=\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\left[\sum_{t=1}^{T}\xi_{t}^{(i,j)}l_{\operatorname{\operatorname{TLS}}}(\tilde{x}_{t}^{(i,j)},y_{i},\theta)+\xi_{T+1}^{(i,j)}l_{\operatorname{\operatorname{TLS}}}(w_{i},y_{i},\theta)\right].$ (13)
where in (12)-(13), optimization over $\nu$ is omitted since the parameter of interest is $\theta$. This loss function is equivalent to minimizing the joint likelihood between $(W,Y)$ conditionally on $X$ in the case of Gaussian distributed errors, i.e. when $\operatorname{\mathbb{F}}_{E}^{0}\equiv N(0,\sigma^{2}_{\epsilon})$ and $\operatorname{\mathbb{F}}_{N}^{0}\equiv N(0,\sigma^{2}_{\nu})$ (see Appendix A.1). The pseudo-code for the Robust-MEM with the TLS objective is outlined in Algorithm 3. We note that similarly to the general NPL framework, Algorithm 3 is parallelizable across the number of Bootstrap iterations. This framework allows for flexible Bayesian parameter inference in MEMs while accounting for prior beliefs on the ME. However, it only allows for inference on Gaussian distributed errors and linear regression functions, which can be limiting and prone to misleading outcomes if the assumptions are misspecified. To address this limitation, we will now introduce a loss function based on the MMD that can accommodate any parametric model and any regression function which might arise in modern statistical and machine learning applications.
input : $(w_{i},y_{i})_{1:n}$, $T$, $B$, $c$, $(\operatorname{\mathbb{F}}_{w_{i}})_{1:n}$
for $j\leftarrow 1$ to $B$ do
       for $i\leftarrow 1$ to $n$ do
      Calculate $\theta^{(j)}=\theta^{*}_{TLS}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ as in (12)
return Posterior bootstrap sample $\theta^{(1:B)}$
Algorithm 3 Robust-MEM with the TLS
@@ 3.4.2 Robust-MEM with the Maximum Mean Discrepancy
The use of distance-based loss functions in the NPL framework was suggested by Dellaporta et al., (2022) to perform robust inference for simulator-based models with the Maximum Mean Discrepancy (MMD). The MMD is defined in a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}_{k}$ with associated kernel $k:\mathcal{Z}\times\mathcal{Z}\rightarrow\mathbb{R}$ and norm $\|\cdot\|_{\mathcal{H}_{k}}$. The MMD has already been used in the context of universal robust regression by Alquier and Gerber, (2024) and is defined through the notion of kernel mean embeddings For a function $f\in\mathcal{H}_{k}$, the kernel mean embedding $\mu_{\operatorname{\mathbb{P}}}\in\mathcal{H}_{k}$ with respect to probability measure $\operatorname{\mathbb{P}}\in\mathcal{P}_{\mathcal{Z}}$ is defined as $\mu_{\operatorname{\mathbb{P}}}(\cdot):=\mathbb{E}_{Z\sim\operatorname{\mathbb{P}}}[k(Z,\cdot)]\in\mathcal{H}_{k}$. The MMD between the probability measures $\operatorname{\mathbb{P}}$ and $\operatorname{\mathbb{Q}}$ is defined as $\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}},\operatorname{\mathbb{Q}})=\|\mu_{\operatorname{\mathbb{P}}}-\mu_{\operatorname{\mathbb{Q}}}\|_{\mathcal{H}_{k}}$. For a parametric model family $\{\operatorname{\mathbb{P}}_{\theta}:\theta\in\Theta\}$ and a sample $\{z_{i}\}_{i=1}^{n}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{P}}^{0}$, the squared MMD can be used as a loss function in (7), resulting in a minimum MMD estimator (Briol et al.,, 2019), given by:
@@
$=\arg\inf_{\theta\in\Theta}\int_{\mathcal{Z}\times\mathcal{Z}}k(z_{1},z_{2})\operatorname{\mathbb{P}}_{\theta}(dz_{1})\operatorname{\mathbb{P}}_{\theta}(dz_{2})-\frac{2}{n}\sum_{i=1}^{n}\int_{\mathcal{Z}}k(z_{i},z_{1})\operatorname{\mathbb{P}}_{\theta}(dz_{1})$
@@
where ^$\hat{\operatorname{\mathbb{P}}}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{z_{i}}$ and $z_{i}:=(x_{i},y_{i})$ for each $i=1,\dots,n$. In practice, this is approximated using a U-statistic and optimized with gradient-based methods. This loss function allows for the specification of nonlinear regression functions $g(\theta,x)$ as well as non-Gaussian distributed errors for both $X$ and $Y$.
@@
We denote by $\operatorname{\mathbb{P}}^{0}$ the joint DGP of $X$ and $Y$, such that $\operatorname{\mathbb{P}}^{0}=\operatorname{\mathbb{P}}^{0}_{X}\operatorname{\mathbb{P}}^{0}_{Y\>|\>\cdot}$. Then the target parameter of interest is:
$\theta^{0}_{\operatorname{\operatorname{MMD}}}$ $:=\arg\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}).$ (14)
In words, the target in (14) represents the parameter value that best explains the relationship between $X$ and $Y$. The uncertainty here is twofold: firstly $\operatorname{\mathbb{P}}^{0}$ is unknown and secondly, we do not have access to samples from $\operatorname{\mathbb{P}}_{X}$, as we only observe realizations $\{(w_{i},y_{i})\}_{i=1}^{n}$ of:
$\operatorname{\mathbb{P}}^{n}(A\times B):=\frac{1}{n}\sum_{i=1}^{n}\delta_{w_{i}}(A)\times\operatorname{\mathbb{P}}_{Y\>|\>X=x_{i}}(B),\quad A\times B\in\mathcal{\sigma}_{\mathcal{Z}}.$
@@
We follow the same nonparametric framework for each of the distributions $\operatorname{\mathbb{P}}_{X\>|\>w_{i}}$ as in (5) and (6). This leads to the NPL posterior density for $\theta$ through the push-forward measure $(\theta^{*}_{\operatorname{\operatorname{MMD}}})^{\#}\left(\mu_{1},\dots,\mu_{n}\right)$ with $\mu_{i}$ as in (9). The Posterior bootstrap algorithm (Algorithm 4) facilitates sampling from this measure by iteratively optimizing the following objective:
@@
$\theta^{*}_{\operatorname{\operatorname{MMD}}}\left(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)}\right)=\arg\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}\left(\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\delta_{y_{i}},\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}\right)$ (15)
@@
where $\operatorname{\mathbb{P}}^{(i,j)}\sim\hat{\mu_{i}}$ for each $i=1,\dots,n$ and if $(X,Y)\sim\operatorname{\mathbb{P}}^{(i,j)}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ then $X\sim\operatorname{\mathbb{P}}^{(i,j)}$ and $Y\>|\>X=x\sim\operatorname{\mathbb{P}}_{g(\theta,x)}$ mirroring the definition of Alquier and Gerber, (2024).
input : $(w_{i},y_{i})_{1:n}$, $T$, $B$, $c$, $(\operatorname{\mathbb{F}}_{w_{i}})_{1:n}$
for $j\leftarrow 1$ to $B$ do
       for $i\leftarrow 1$ to $n$ do
      Obtain $\theta^{(j)}=\theta^{\ast}_{\text{MMD}}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ as in 15 using numerical optimization.
return Posterior bootstrap sample $\theta^{(1:B)}$
Algorithm 4 Robust-MEM with the MMD
4 Theoretical results for Robust-MEM with the MMD
@@
In this section, we provide generalization error bounds for the Robust-MEM framework with the MMD. We define the target distribution for the Berkson and Classical cases and outline technical assumptions in Section 4.1. The main results are presented in Section 4.2. Finally, we analyze the theoretical results under common assumptions in the ME literature, including known error distribution and Gaussian-distributed ME, in Section 4.3.
@@ 4.1 Target distribution & assumptions
The goal is to bound the expected error in terms of the MMD between the true joint distribution of $X,Y$, namely $\operatorname{\mathbb{P}}^{0}$, and the estimated model. The Berkson and Classical error cases lead to different forms of $\operatorname{\mathbb{P}}^{0}$, due to the dependence structures of $X$, $N$, $W$.
@@
$\operatorname{\mathbb{P}}^{0}(W\in dw,X\in dx,Y\in dy)=\frac{1}{n}\sum_{i=1}^{n}\delta_{w_{i}}(dw)\operatorname{\mathbb{P}}^{0}_{X\>|\>w}(dx)\operatorname{\mathbb{P}}^{0}_{Y\>|\>x}(dy)$ (16)
where $\operatorname{\mathbb{P}}^{0}_{X\>|\>w}$ and $\operatorname{\mathbb{P}}^{0}_{Y\>|\>x}$ denote the true conditional distributions of $X\>|\>W=w$ and $Y\>|\>X=x$ respectively. By marginalizing out $W$ we obtain the target joint distribution of $X,Y$ which is equal to:
[…]
@@
$\operatorname{\mathbb{P}}^{0}(X\in dx,W\in dw,Y\in dy)=\operatorname{\mathbb{P}}^{0}_{X}(dx)\operatorname{\mathbb{P}}^{0}_{W\>|\>x}(dw)\operatorname{\mathbb{P}}^{0}_{Y\>|\>x}(dy).$ (19)
whereas the target joint distribution of $X,Y$ is simply
[…]
We can now define the desired generalization error given the target joint and marginal distributions. For random samples $(\operatorname{\mathbb{P}}^{i}\sim\mu_{i})_{1:n}$, with $\mu_{i}$ as in (9) define $\operatorname{\mathbb{P}}:=(\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n})$. The estimated model $\operatorname{\mathbb{P}}_{\theta^{*}(\operatorname{\mathbb{P}})}$, relative to the noise-free covariates, for each of the Berkson and Classical ME cases is:
[…]
respectively. The goal is to examine how well $\operatorname{\mathbb{P}}^{0}_{\theta^{*}(\operatorname{\mathbb{P}})}$ approximates $\operatorname{\mathbb{P}}^{0}$ in expectation if we had access to the distribution of the true covariates. That is, we check if the resulting model captures the relationship between the true covariates and the response variable.
In what follows, we provide results for the exact DP case for clarity. The exact DP is represented through the stick-breaking process (Sethuraman,, 1994) as follows: for each $\operatorname{\mathbb{P}}^{i}\sim\mu_{i}:=\operatorname{\operatorname{DP}}(c+1,\operatorname{\mathbb{F}}_{w_{i}}^{\prime})$:
@@
Using this notation we also define the joint model under samples from the DP posteriors:
$\operatorname{\mathbb{P}}_{\theta}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{i}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{\infty}\xi_{k}^{i}\delta_{\tilde{x}_{k}^{i}}\operatorname{\mathbb{P}}_{g(\theta,\tilde{x}_{k}^{i})}.$ (24)
Recall that $\mathcal{X}$ and $\mathcal{Y}$ are two topological spaces, equipped respectively with the Borel $\sigma$-algebra $\mathfrak{S}_{\mathcal{X}}$ and $\mathfrak{S}_{\mathcal{Y}}$, and let $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, $\mathfrak{S}_{\mathcal{Z}}=\mathfrak{S}_{\mathcal{X}}\otimes\mathfrak{S}_{\mathcal{Y}}$ and $\mathcal{P}(\mathcal{Z})$ be the set of probability distributions on $(\mathcal{Z},\mathfrak{S}_{\mathcal{Z}})$. We let $k:\mathcal{Z}^{2}\rightarrow\mathbb{R}$ be a reproducing kernel on $\mathcal{Z}$ and denote by $\mathcal{H}_{k}$ the reproducing kernel Hilbert space (RKHS) over $\mathcal{Z}$ with $k$. We assume throughout this work that the following standard condition on $k$ holds:

### v3

@@ 3.3 Prior Elicitation
We now focus on the concentration parameter of the DP prior in (3). It is clear from the posterior in (4) that $c=0$ corresponds to the prior belief that there is no ME present on the covariates resulting in all the mass of the posterior centering measure concentrating at $\delta_{w_{i}}$. This is in contrast to the standard NPL setting (Sec. 2) where $c=0$ corresponds to a noninformative prior. Moreover, it is evident that $c=1$ results in splitting the weight of the DP posterior centering measure $\operatorname{\mathbb{F}}_{w_{i}}^{\prime}$ equally between the prior centering measure $\operatorname{\mathbb{F}}_{w_{i}}$ and the Dirac measure on the observation $w_{i}$. Lastly, when $c\rightarrow\infty$ we see that all the weight is placed in the prior belief. The interplay between parameter $c$, ME variance, and prior specification can be further demonstrated with an example of polynomial regression with Berkson, normally distributed ME in Fig. 3.
Figure 3: Average MSE in polynomial regression with Berkson, Gaussian ME over 50 replications. Left: Robust-MEM method with a well-specified prior. Right: Robust-MEM method with a misspecified (Student-t with 3 degrees of freedom) prior.
Here, the average MSE over the three parameters of interest is reported for the three limiting choices of $c$, for an increasing ME variance. In the case of a well-specified prior (left), when the ME variance is close to zero, i.e. when there is no ME, all three $c$ values perform similarly since the prior correctly indicates no ME, making inference data-driven. The choices of $c>0$ perform similarly, yielding a lower MSE for nonzero ME variance. Conversely, in the case where the true ME distribution is Gaussian and the prior centering measure $\operatorname{\mathbb{F}}_{N}^{0}$ is a Student-t distribution with 3 degrees of freedom (right plot), $c=1$ corresponds to the lowest average MSE, as expected. A setting of $c=1$ appropriately balances prior beliefs and observations, while $c=0$ wrongly assumes no ME, and $c=100$ overly weights the misspecified prior. In general, overlapping standard deviations in most cases indicate the robustness of the method to $c$ values. In the absence of strong prior beliefs on the ME distribution, setting $c=1$ splits the weight sensibly between prior beliefs and observations.
@@ 3.4 Choice of Loss Function
In this section, we explore the adaptability of the Robust-MEM framework to various prevalent cases of ME. We propose two distinct loss functions suitable to address these scenarios comprehensively. The initial choice stems from the TLS objective, offering rapid and robust solutions, particularly tailored for scenarios such as linear regression and Gaussian-distributed ME. To ensure broad applicability across diverse model families and regression relationships, we introduce the MMD-based loss function as a further approach.
@@ 3.4.1 Robust-MEM with the Total Least Squares
We first suggest a loss function based on the TLS objective function. Consider a linear regression model. Then we can define the loss function $l_{\operatorname{\operatorname{TLS}}}:\mathcal{Z}\times\Theta\rightarrow\mathbb{R}$ as $l_{\operatorname{\operatorname{TLS}}}(x_{i},y_{i},\theta):=\|\nu_{i}\|_{2}^{2}+(y_{i}-\theta^{T}(x_{i}+\nu_{i}))^{2}.$ Using this loss function in (7), the minimization objective for each set of DP samples $(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ as in (8) is
$\theta^{\star}_{TLS}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$
$\quad\quad=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{(x,y)\sim\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\times\delta_{y_{i}}}\left[l_{\operatorname{\operatorname{TLS}}}(x,y,\theta)\right]$
$\quad\quad=\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\left[\sum_{t=1}^{T}\xi_{t}^{(i,j)}l_{\operatorname{\operatorname{TLS}}}(\tilde{x}_{t}^{(i,j)},y_{i},\theta)+\xi_{T+1}^{(i,j)}l_{\operatorname{\operatorname{TLS}}}(w_{i},y_{i},\theta)\right]$ (10)
where in (10), optimisation over $\nu$ is omitted since the parameter of interest is $\theta$. This loss function is equivalent to minimizing the joint likelihood between $(W,Y)$ conditionally on $X$ in the case of Gaussian distributed errors, i.e. when $\operatorname{\mathbb{F}}_{E}^{0}\equiv N(0,\sigma^{2}_{\epsilon})$ and $\operatorname{\mathbb{F}}_{N}^{0}\equiv N(0,\sigma^{2}_{\nu})$ (see Appendix C.1). The pseudo-code for the Robust-MEM with the TLS objective is outlined in Algorithm 3 (Appendix E). This framework allows for flexible Bayesian parameter inference in MEMs while accounting for prior beliefs on the ME. However, it only allows for inference on Gaussian distributed errors and linear regression functions, which can be limiting and prone to misleading outcomes if the assumptions are misspecified. To address this limitation, we introduce a loss function based on the MMD.
@@ 3.4.2 Robust-MEM with the Maximum Mean Discrepancy
[…]
@@
$=\arg\inf_{\theta\in\Theta}\int_{\mathcal{Z}\times\mathcal{Z}}k(z_{1},z_{2})\operatorname{\mathbb{P}}_{\theta}(dz_{1})\operatorname{\mathbb{P}}_{\theta}(dz_{2})-\frac{2}{n}\sum_{i=1}^{n}\int_{\mathcal{Z}}k(z_{i},z_{1})\operatorname{\mathbb{P}}_{\theta}(dz_{1}).$
@@
In practice, this is approximated using a U-statistic and optimized with gradient-based methods. This loss function is suitable for general regression models and hence allows for the specification of nonlinear regression functions $g(\theta,x)$ as well as non-Gaussian distributed errors for both $X$ and $Y$.
@@
Recall that $\operatorname{\mathbb{P}}^{0}$ denotes the joint DGP of $X$ and $Y$, such that $\operatorname{\mathbb{P}}^{0}=\operatorname{\mathbb{P}}^{0}_{X}\times\operatorname{\mathbb{P}}^{0}_{Y\>|\>X}$. Then the target parameter of interest is:
$\theta^{0}_{\operatorname{\operatorname{MMD}}}:=\arg\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}),$
@@
representing the parameter value that best explains the relationship between $X$ and $Y$. The uncertainty is twofold: firstly $\operatorname{\mathbb{P}}^{0}$ is unknown and secondly, we do not have access to samples from $\operatorname{\mathbb{P}}_{X}$, as we only observe realizations $\{(w_{i},y_{i})\}_{i=1}^{n}$ of: $\operatorname{\mathbb{P}}^{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{w_{i}}\times\operatorname{\mathbb{P}}_{Y\>|\>w_{i}}.$ We follow the same nonparametric framework for each of the distributions $\operatorname{\mathbb{P}}_{X\>|\>w_{i}}$ as in (3) and (4). This leads to the NPL posterior density for $\theta$ through the push-forward measure $(\theta^{\star}_{\operatorname{\operatorname{MMD}}})^{\#}\left(\mu_{1},\dots,\mu_{n}\right)$ with $\mu_{i}$ as in (6). The Posterior bootstrap algorithm (Algorithm 4) facilitates sampling from this measure by iteratively optimizing the following objective:
$\theta^{\star}_{\operatorname{\operatorname{MMD}}}\left(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)}\right)$
@@
$\quad\quad=\arg\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}\left(\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\times\delta_{y_{i}},\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}\right)$ (11)
@@
where $\operatorname{\mathbb{P}}^{(i,j)}\sim\hat{\mu_{i}}$ for each $i=1,\dots,n$ and if $(X,Y)\sim\operatorname{\mathbb{P}}^{(i,j)}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ then $X\sim\operatorname{\mathbb{P}}^{(i,j)}$ and $Y\>|\>X=x\sim\operatorname{\mathbb{P}}_{g(\theta,x)}$ mirroring the definition in Alquier and Gerber, (2024).
4 Generalization Error of Robust-MEM with the MMD
@@
In this section, we present generalisation error bounds for the Robust-MEM framework with the MMD. We define the target distribution for the Berkson and Classical cases and outline key technical assumptions in section 4.1. The main results are presented in section 4.2. Finally, we analyze the theoretical results under common ME assumptions, such as known error distribution and Gaussian-distributed ME, in Section 4.3. Before we start the theretical assesment we summarise the notation used for each error structure in Table 1.
Table 1: Summary of notation for Berkson and Classical error models.
Berkson Classical
Target distribution of $(X,Y)$ $\operatorname{\mathbb{P}}^{0}:=\operatorname{\mathbb{P}}^{0,n}_{X}\times\operatorname{\mathbb{P}}^{0}_{Y\>|\>X}$ $\operatorname{\mathbb{P}}^{0}:=\operatorname{\mathbb{P}}^{0}_{X}\times\operatorname{\mathbb{P}}^{0}_{Y\>|\>X}$
Joint model of $(X,Y)$ $\operatorname{\mathbb{P}}^{0}_{\theta}:=\operatorname{\mathbb{P}}^{0,n}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ $\operatorname{\mathbb{P}}^{0}_{\theta}:=\operatorname{\mathbb{P}}^{0}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$
@@ 4.1 Target Distribution & Assumptions
The goal is to bound the expected error in terms of the MMD between the true joint distribution of $X,Y$, namely $\operatorname{\mathbb{P}}^{0}$, and the estimated model. The Berkson and Classical error cases lead to different forms of $\operatorname{\mathbb{P}}^{0}$, due to the dependence structures of $X$, NN, $W$.
@@
$\operatorname{\mathbb{P}}^{0}(X\in dx,Y\in dy)$ $:=\operatorname{\mathbb{P}}^{0,n}_{X}(dx)\times\operatorname{\mathbb{P}}^{0}_{Y\>|\>X}(dy),$ (12)
@@
$\operatorname{\mathbb{P}}^{0}(X\in dx,Y\in dy)=\operatorname{\mathbb{P}}^{0}_{X}(dx)\times\operatorname{\mathbb{P}}^{0}_{Y\>|\>X}(dy).$ (14)
For random samples $\{\operatorname{\mathbb{P}}^{i}\sim\mu_{i}\}_{i=1}^{n}$, with $\mu_{i}$ as in (6) define $\operatorname{\mathbb{P}}:=(\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n})$. The estimated model $\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}$, relative to the noise-free covariates, for the Berkson and Classical cases is:
$\text{\tt(Berkson)}\quad\operatorname{\mathbb{P}}^{0}_{\theta^{\star}(\operatorname{\mathbb{P}})}(dx\times dy)$ $:=\operatorname{\mathbb{P}}^{0,n}_{X}(dx)\times\operatorname{\mathbb{P}}_{g(\theta^{\star}(\operatorname{\mathbb{P}}),\cdot)}(dy)\quad\text{or}$ (15)
$\text{\tt(Classical)}\quad\operatorname{\mathbb{P}}^{0}_{\theta^{\star}(\operatorname{\mathbb{P}})}(dx\times dy)$ $:=\operatorname{\mathbb{P}}^{0}_{X}(dx)\times\operatorname{\mathbb{P}}_{g(\theta^{\star}(\operatorname{\mathbb{P}}),\cdot)}(dy)$ (16)
The goal is to examine how well $\operatorname{\mathbb{P}}^{0}_{\theta^{\star}(\operatorname{\mathbb{P}})}$ approximates $\operatorname{\mathbb{P}}^{0}$ in expectation if we had access to the distribution of the true covariates. That is, we check if the resulting model captures the relationship between the true covariates and the response variable.
In what follows, we consider the exact DP case for clarity. This is represented through the stick-breaking process (Sethuraman,, 1994) such that for each $\operatorname{\mathbb{P}}^{i}\sim\mu_{i}:=\operatorname{\operatorname{DP}}(c+1,\operatorname{\mathbb{F}}_{w_{i}}^{\prime})$:
@@
Using this notation, we also define the joint model under samples from the DP posteriors:
$\operatorname{\mathbb{P}}_{\theta}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{i}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{\infty}\xi_{k}^{i}\delta_{\tilde{x}_{k}^{i}}\times\operatorname{\mathbb{P}}_{g(\theta,\tilde{x}_{k}^{i})}.$ (18)
Recall that $\mathcal{X}$ and $\mathcal{Y}$ are two topological spaces, equipped respectively with the Borel $\sigma$-algebra $\mathfrak{S}_{\mathcal{X}}$ and $\mathfrak{S}_{\mathcal{Y}}$, and let $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, $\mathfrak{S}_{\mathcal{Z}}=\mathfrak{S}_{\mathcal{X}}\otimes\mathfrak{S}_{\mathcal{Y}}$ and $\mathcal{P}_{\mathcal{Z}}$ be the set of probability distributions on $(\mathcal{Z},\mathfrak{S}_{\mathcal{Z}})$. We let $k:\mathcal{Z}^{2}\rightarrow\mathbb{R}$ be a reproducing kernel on $\mathcal{Z}$ and denote by $\mathcal{H}_{k}$ the reproducing kernel Hilbert space (RKHS) over $\mathcal{Z}$ with $k$. Examples of kernels include the Radial Basis Function (RBF) or Gaussian kernel defined as $k(x,y):=\exp\left\{-\frac{1}{2l^{2}}\|x-y\|^{2}\right\}$ for lengthscale $l$, the polynomial kernel of degree $d$ defined as $k(x,y):=(\left<x,y\right>+c)^{d}$ for some constant $c$ and the exponential kernel defined as $k(x,y):=\exp\{\left<x,y\right$>\}. Apart from the reproducing property, there are two other useful properties of kernels that we use in this paper. First, characteristic kernels are kernels that induce an injection between probability measures and kernel mean embeddings, more formally:
Definition 4.0.1 (Characteristic kernel. Sriperumbudur et al., (2011)).
A bounded and measurable kernel $k$ is called characteristic if the map $f:\mathcal{P}_{\mathcal{Z}}\rightarrow\mathcal{H}_{k}$ defined as $f(\operatorname{\mathbb{P}})=\mu_{\operatorname{\mathbb{P}}}$ is an injection.
This property guarantees that the MMD is a divergence, meaning that
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}},\operatorname{\mathbb{Q}})=0\iff\operatorname{\mathbb{P}}\equiv\operatorname{\mathbb{Q}}.$
Another useful property that a kernel may possess is translation invariance. This occurs when a linear perturbation of both inputs does not change the kernel value.
Definition 4.0.2 (Translation invariant kernel.).
The kernel $k$ is translation invariant if there exists a bounded, continuous, real-valued, positive definite function $\psi$ on $\mathcal{X}\subseteq\operatorname{\mathbb{R}}^{d}$ such that $k(x,y)=\psi(x-y)$ for all $x,y\in\mathcal{X}$.
We assume throughout this work that the following standard condition on $k$ holds: